OpenStudy (anonymous):

Use the linearization (1+x)^k=1+kx to approximate 1.002^100

OpenStudy (anonymous):

1+(0.002)100 =1.2

OpenStudy (turingtest):

(1.002)=(1+0.002) so x=0.002 linearization: 1+(0.002)100=1.2

OpenStudy (anonymous):

Thank you! So x is the base and k is the exponent?

OpenStudy (turingtest):

x+1 is the base and k is the exponent in the original equation, but when it is linearized k becomes a coefficient.

OpenStudy (anonymous):

Okay okay I see now. Thank you